# Contrastive Preference Optimization: Pushing the Boundaries of LLM Performance in Machine Translation

### 18 Jan 2024 v2

@@ Appendix B Prompts for Translations
Appendix C Simplifying The Re-Written DPO Loss Function
In this section, we provide a detailed explanation of the approximation used to optimize the loss function from Equation 3 to Equation 6. We begin by expanding the Sigmoid function in Equation 3:
Given that is a fixed model and does not participate in gradient calculations or parameter updates, the above loss function can be simplified by omitting the term :
@@ D.2 Influence on Performance
Results and Analysis We show the detailed results for enxx and xxen in Table 12 and 13, respectively. The inclusion of human-labeled preference data does not significantly enhance overall translation performance. For enzh, marginal improvements are observed, though they are minimal. Conversely, for ende, a slight decline in performance is noted. In summary, the addition of human-labeled data shows no substantial difference in the enxx direction, and a minor decrease in performance for xxen on average. We hypothesize that the limited impact of these human-labeled data may stem from a high proportion of tied evaluations and potential human bias in the evaluation process. For instance, there are instances where the author consider GPT-4’s translations to be superior, while human evaluators favor those produced by Google Translate.
@@ E.2 Systems For WMT’23
Appendix F Full Results of WMT’23

### 02 Feb 2024 v3

@@ Appendix B Prompts for Translations
Appendix C Theory
C.1 Proof of The Upper Boundary
Theorem 1.
When is defined as , an ideal policy that precisely aligns with the true data distribution of preferred data, the DPO loss is upper bounded by , where is a constant.
Proof. represents an ideal policy that perfectly aligns the true data distribution of the preferred data. Hence, for any given data point from the preference dataset , the conditions and hold true. Consequently, under this setup, the predictions for preferred data do not require reweighting by the reference model, and the DPO loss can be reformulated as follows :
After expanding the Sigmoid function, the loss becomes to:
Given that is a fixed model and does not participate in gradient calculations or parameter updates, the above loss function is equivalent when we omit the term . Therefore, optimizing is equivalent to optimizing as we define below:
@@ C.1 Proof of The Upper Boundary
Therefore, is upper bounded by , where .
C.2 BC Regularizer Simplification
The contrastive preference optimization is originally defined as minimizing under the constraint of minimizing the difference between preferred data distribution and outputs of the learnable policy:
This is equivalent to the following objective via Lagrangian duality:
where is a hyperparamter and we set to 1. The optimization can be further optimized by expanding the KL divergence:
This results in the final formulation of our CPO loss function.
@@ D.2 Influence on Performance
Results and Analysis We show the detailed results for enxx and xxen in Table 10 and 11, respectively. The inclusion of human-labeled preference data does not significantly enhance overall translation performance. For enzh, marginal improvements are observed, though they are minimal. Conversely, for ende, a slight decline in performance is noted. In summary, the addition of human-labeled data shows no substantial difference in the enxx direction, and a minor decrease in performance for xxen on average. We hypothesize that the limited impact of these human-labeled data may stem from a high proportion of tied evaluations and potential human bias in the evaluation process. For instance, there are instances where the author consider GPT-4’s translations to be superior, while human evaluators favor those produced by Google Translate.
@@ E.2 Systems For WMT’23
Appendix F Estimated Accuracy with Human Agreements
In the paper, we adopt a new approach for highlighting improvements within tables, moving beyond the standard practice of specifying a static improvement threshold in metric by score . Instead, our threshold is dynamic, calibrated to the minimal metric difference in metric that yields a perceptible distinction between two systems as recognized by humans (Kocmi et al., 2024). For instance, to align with human judgments at an 80% concordance rate, the required improvement margin is for both KIWI-XXL and COMET-XXL, and for KIWI-22. A comprehensive delineation of these thresholds can be found in Table 13.
| Estimated Accuracy | Coin toss 50% | 55% | 60% | 65% | 70% | 75% | 80% | 85% | 90% | 95% |
| BLEU | 0.27 | 0.52 | 0.78 | 1.06 | 1.39 | 1.79 | 2.34 | 3.35 | - | - |
| Comet-22 | 0.03 | 0.10 | 0.18 | 0.26 | 0.35 | 0.45 | 0.56 | 0.71 | 0.94 | 1.53 |
| KIWI-22 | 0.01 | 0.08 | 0.16 | 0.24 | 0.33 | 0.42 | 0.53 | 0.67 | 0.85 | 1.18 |
| XCOMET-XXL | 0.02 | 0.19 | 0.37 | 0.56 | 0.76 | 0.98 | 1.24 | 1.55 | 1.99 | 2.74 |
| KIWI-XXL | 0.06 | 0.22 | 0.39 | 0.57 | 0.77 | 0.98 | 1.24 | 1.58 | 2.08 | 3.39 |
Appendix G Full Results of WMT’23
@@
Appendix H Evaluation on ALMA-R with Non-Comet Metric
Concerns may arise regarding the similar training procedure of COMET metrics, leading to high correlation among COMET models, which potentially undermine the validity of our analysis in Section 5.1. To address this, we also consider BLEURT-20 (Sellam et al., 2020), a non-COMET and neural-based (but reference-based evaluation) metric. We present BLEURT scores for ALMA-13B-LoRA and ALMA-13B-R in Table 15. Notably, even when preference data is constructed using COMET-based evaluations, significant improvements in non-COMET scores are observed. This strengthens our findings that translations produced by ALMA-R are indeed superior and robust.
| BLEURT-20 | de | cs | is | zh | ru | Avg. |
| Translating to English (xxen) | ||||||
| ALMA-13B-LoRA | 73.20 | 76.65 | 75.87 | 67.37 | 76.7 | 73.96 |
| ALMA-13B-R | 73.62 | 76.94 | 76.98 | 69.48 | 76.91 | 74.79 |
| Translating from English (enxx) | ||||||
| ALMA-13B-LoRA | 75.51 | 80.93 | 73.19 | 70.54 | 74.94 | 75.02 |
| ALMA-13B-R | 77.20 | 81.87 | 73.43 | 71.51 | 76.19 | 76.04 |
Appendix I The Effectiveness of The BC Regularizer for DPO
The DPO loss can also be utilized by adding our additional BC regularizer:
In Table 16, we demonstrate that incorporating into the DPO objective results in notable enhancements for translations both to and from English. This observation hints at why , as an approximation of , performs effectively, while the original DPO loss does not. It appears that the DPO loss lacks the BC regularizer, which steers the model towards the preferred data distribution. Although combining DPO with the BC regularizer could yield similar performance to CPO, it incurs double the memory cost and FLOPs per token in the forward pass. The original DPO loss shows the possibility of failure to improve the model performance in preference learning, so we here highlight the significance of incorporating BC regularization. Importantly, Table 16 shows that is a successful approximation of the DPO loss, offering savings in memory and speed, and it can even outperform the original BC-regularized DPO loss .
| Loss Objective | KIWI-22 | KIWI-XXL | XCOMET | Memory Cost | FLOPs/tok |
| Translating to English (xxen) | |||||
| 80.51 | 81.36 | 86.58 | 2 | 2 | |
| 81.28 | 82.42 | 89.05 | 2 | 2 | |
| (CPO) | 81.33 | 82.43 | 89.11 | 1 | 1 |
| Translating from English (enxx) | |||||
| 82.27 | 82.07 | 92.25 | 2 | 2 | |
| 83.13 | 84.74 | 93.53 | 2 | 2 | |
| (CPO) | 83.34 | 85.74 | 94.05 | 1 | 1 |